OpenStudy (anonymous):

For the equation x^2+y^2-8x-2y-19=0 do the following. a. find the center (h,k) and the radius r of the circle. b. graph the circle. c. Find the intercepts, if any.

OpenStudy (anonymous):

you have to complete the square twice do you know how to do that?

OpenStudy (anonymous):

Yes, I do! I just don't really know how to do that with this problem...

OpenStudy (anonymous):

\[x^2-8x+y^2-2y=19\]

OpenStudy (anonymous):

half of 8 is 4, half of 2 is 1 go right to \[(x-4)^2+(y-1)^2=19+4^2+1^2\] and work from there

OpenStudy (anonymous):

ok! Thank you! I have to write that down

OpenStudy (anonymous):

yes you should write it down yw

OpenStudy (anonymous):

I meant to write it down so i can work it out and transfer it on to my online class... :)